# Persona-driven Simulation of Voting Behavior in the European Parliament with Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) display remarkable capabilities to understand or even produce political discourse, but have been found to consistently display a progressive left-leaning bias. At the same time, so-called persona or identity prompts have been shown to produce LLM behavior that aligns with socioeconomic groups that the base model is not aligned with. In this work, we analyze whether zero-shot persona prompting with limited information can accurately predict individual voting decisions and, by aggregation, accurately predict positions of European groups on a diverse set of policies. We evaluate if predictions are stable towards counterfactual arguments, different persona prompts and generation methods. Finally, we find that we can simulate voting behavior of Members of the European Parliament reasonably well with a weighted F1 score of approximately 0.793. Our persona dataset of politicians in the 2024 European Parliament and our code are available at https://anonymous.4open.science/r/european_parliament_simulation-7285.
@@
Maximilian Kreutner1, Marlene Lutz1, Markus Strohmaier1,2 1University of Mannheim 2GESIS - Leibnitz Institute for the Social Sciences, Cologne Correspondence: maximilian.kreutner@uni-mannheim.de; marlene.lutz@uni-mannheim.de; markus.strohmaier@uni-mannheim.de
@@ 1 Introduction
Large Language Models (LLMs) are trained on vast amounts of diverse data, encompassing a wide range of attitudes, opinions, and values. Given their broad exposure, it remains unclear to what extent these models inherit various perspectives and how these perspectives might resurface in their responses. As a result, there is a growing body of research dedicated to investigating the latent viewpoints of LLMs, particularly in relation to their political leaning feng2023pretraining; wright-etal-2024-llm; rottger2024political. Studies have found that, by default, LLMs tend to exhibit progressive, left-center political positions rutinowski2024selfperception; taubenfeld2024systematic; hartmann2023political; socsci12030148. At the same time, researchers have demonstrated that when properly adapted (e.g., by fine-tuning or prompting), LLMs can effectively understand bernardelle2025mapping; heseltine2024large; wu2023large and reproduce bachmann2025adaptive; chalkidis2024llamaEU the positions of various political groups.
Research objective: Building on these early demonstrations, our work systematically explores to what extent LLMs can represent the political votes of individual politicians. Specifically, we aim to simulate the voting behavior of all Members of the European Parliament (MEPs) in 2024, by leveraging the ability of LLMs to role-play as different personas. This type of persona prompting has been shown to allow LLMs to make predictions that better align with the points of view of various sociodemographic groups in certain contexts argyle2023onemany; tseng2024twopersona.
The case of the European Parliament is particularly interesting, as MEPs are often part of both a national party and a European group. Therefore, MEPs’ personal preferences and national party lines can significantly influence the decision to dissent from their corresponding European Group willumsen2023policy. Although individual voting decisions in the early 2000s were still mostly decided by party line of national parties hix2002parliamentary, the European Parliament has since gone through a phase of "Empowerment", where majority coalitions are policy-specific, which is unique compared to most national parliaments hix2013empowerment.
@@
Approach: We use personas to prompt four LLMs to adopt the identity of each MEP and investigate the extent to which these models can simulate their voting behavior (cf. Figure 1). This includes predicting individual decision-making of MEPs, who often balance competing interests, and the group line on a given policy. We compare different models, persona prompts and generation methods. We perform additional ablation studies to find out which attributes are the best predictor for LLMs.
Results: We find that in the majority of cases persona-driven LLMs can accurately predict individual votes, achieving a weighted F1 score of 0.793, but do not predict abstentions well. Additionally, persona-driven LLMs are able to correctly predict group positions on a diverse set of topics while also diverging from it at a rate comparable to real-world politicians. Our ablation study supports current political research, as the national party is the strongest predictor of the final vote.
In persona-driven simulations, we find evidence supporting three key observations. First, more detailed personas can help weaker models generate more accurate predictions. Second, prompting models with open-text reasoning before forcing multiple-choice options steer models more. Third, models struggle to convincingly replicate views of personas that diverge from their inherent biases.
@@ 2 Related Work
Predicting politicians’ voting behavior is an active research area in both political science and NLP. Early work used word frequency, TF-IDF, fine-tuned embeddings, CNNs, or LSTMs to analyze bills and politician profiles gerrish2011predicting; kraft2016embedding; nay2017predicting; bonica2018inferring; kornilova2018party; yang2021joint. Recent methods incorporate social media and socioeconomic data with graph-based models for richer context karimi2020multi; mou2021align; feng2022par; mou2023uppam. Agent Based approaches with Large Language Models were first utilized by li2025political and have shown to outperform previous methods. However, existing studies focus solely on the simpler two-party U.S. system. For the European Parliament only traditional Machine Learning methods e.g. Random Forest have been tested and have shown to perform only slightly better than the majority baseline guadarrama2025knowing. In addition, most recent approaches depend on access to vast amount of past voting information or social media data for each politician, which is not always feasible in real-world scenarios. In contrast, our approach requires only publicly available information about politicians and any form of debate content.
Previous research has consistently shown that various LLMs have a bias toward progressive left-of-center political views, by evaluating them on survey questions or different downstream tasks chalkidis2024llamaEU; chalkidis2024investigating; taubenfeld2024systematic; rutinowski2024selfperception; batzner2024germanpartiesqa; feng2023pretraining; hartmann2023political. Furthermore, human feedback-tuned LLMs show an even larger bias compared to base models potter2024hidden; santurkar2023opinions.
Several studies have demonstrated that LLMs can engage in "role-playing" by adopting specific personas within their contextual framework tseng2024twopersona; chen2024persona. In fact, they are so proficient at this that responses generated while impersonating politicians and other public figures are often perceived as more authentic than those given by the individuals themselves herbold2024large. This role-playing capability has numerous applications. For instance, by using "silicon samples", personas modeled with various socioeconomic attributes, LLMs can effectively replicate the viewpoints of demographically diverse subpopulations or predict election and voting behavior ma2024algorithmic; heyde2024voxpopuli; argyle2023onemany; yu2024large; zhang2024electionsim; bradshaw2024llm; yang2024llm; jiang2024donald. Additionally, generating agents based on long form interviews enables LLMs to reproduce participants’ responses on the General Social Survey with $85\%$ accuracy, compared to their own answers given two weeks later park2024generative. We aim to use this capability to predict voting behavior across European groups, particularly those with whom LLMs are less aligned with.
@@ 3 Data Collection
We collect vote information from HowTheyVote [^1] https://howtheyvote.eu/, a project dedicated to making European Parliament results transparently accessible to the public. Our focus is on roll-call votes, the only type of vote in the European Parliament where the individual votes of all MEPs are recorded and published. Of all 1,688 roll-call votes that were held in the ninth European parliament in the year 2024, we further filter for votes that are associated with a press release and a debate, resulting in a total of 47 votes displayed in Table 5 in Appendix D. Out of the 47 votes, 44 votes are proposals or amendments to proposals that are initiated by the European Commission, while the remaining three votes are resolutions initiated by the European Parliament itself. For the sake of simplicity, from this point on we will refer to all of these votes as proposals. Each MEP can vote in favor of (FOR), against (AGAINST), abstain (ABSTENTION) from a proposal, or not vote at all. While active abstentions are often strategic decisions muhlbock2017legislators, the reasons for not voting are not recorded and can vary, such as illnesses or other responsibilities. As a result, we focus solely on predicting votes categorized as FOR, AGAINST, or ABSTENTION.
@@ 4.1 Models
To prevent data leakage, we choose a model with a training cut-off prior to the earliest vote in our dataset on 16.01.2024: Llama3 dubey2024llama. In addition, as Language Models can show bias towards positions of their creators buyl2024large, we choose a second model that was not developed in the west with comparable performance on the Open LLM Leaderboard openllmleaderboardv2: Qwen2.5555Qwen2.5’s pretraining cutoff date is not officially disclosed. We considered the risk of data leakage but found it unlikely to be a confounding factor as the results are comparably worse to LLama3. yang2024qwen2. More specifically, we use the models:
@@
Additionally, we experimented with Mixtral jiang2024mixtral as a European model, but observed that it did not adhere well to our prompt format, as described in Appendix F. Therefore, we focus on LLama and Qwen for the main section of our paper.
@@ 4.2 Prompt Construction
Persona Construction. To create persona descriptions for each MEP, we consider two strategies. First, we create a combination of demographic and political affiliation attributes, including the full name, gender, age at the time of voting, birthplace, the country they represent, the European group and the national party they are affiliated with at the time of voting. This approach allows us to precisely control the persona information that the LLM is provided with and enables us to estimate the influence of individual attributes on voting behavior. Second, we use Llama3-70B to summarize the English Wikipedia article associated with each MEP. We investigate whether providing the LLM with longer contextual persona information can improve voting accuracy, similar to the approach used by park2024generative who used long form interview transcripts to simulate personas.
Proposal Information. It is challenging to provide a short and unbiased description of proposals. The name of the proposal often often lacks sufficient detail; for example, one proposal in our dataset is simply titled "Driving licenses". We considered using the official document for each proposal but found that the consolidated legislative document is too long for the context window of LLMs. On the other hand, official press releases, news reports, and summaries provided by the European Committee typically focus only on the positive aspects of the proposals, which could potentially bias the LLMs to always vote in favor of the proposals. Instead, we adopt the theory that a common group position is established by the group’s policy experts and then later adopted by other MEPs ringe2009decides. We choose to use the speeches of each group representative, who are allocated speaking time before each vote to present the group’s position on the respective proposal. The speeches are presented to the LLM in randomized order, without disclosing the speaker’s identity or group affiliation. Names of politicians, groups and parties appearing in speeches were anonymized by substituting them with placeholders (e.g. [GROUP]; cf. Appendix E.1). We aim to provide a balanced overview of the pro- and contra-arguments regarding the implementation of the proposal, while avoiding directly revealing each group’s position directly. We note that on average, each speech is nearly 1,500 characters long and makes up the majority of the prompt.
Response Format. We instruct the model to respond with exactly one of the options FOR, AGAINST or ABSTENTION. We further compare two different prompting strategies; reasoning (r), where the model responds first with an open text reasoning chain before choosing one of the options and no reasoning (nr), where the model answers with one of the options directly. We prompt the model three times for each persona. In all runs, we use a temperature of $0.6$. Our full prompts can be found in Appendix G.
@@ 5 Results
Since the distribution of ground truth votes is heavily imbalanced, with approximately $77\%$ FOR votes, $17\%$ AGAINST votes and $6\%$ ABSTENTION votes, we decide for the mean weighted F1 score across three runs as evaluation measure. We report that the responses of all models are robust, i.e. in $87.1\%$ of cases, the models predict the same vote for each persona across all three runs, leading to low standard deviation of <0.002<0.002 across all runs. We analyze performance along three categories:
@@ 5.1 Persona Results
We show our main results in Table 1. The highest weighted F1-score of $0.793$ is achieved by Llama3-70B when employing reasoning with attribute-based persona prompts. Consistent with expectations, larger models surpass their smaller counterparts in performance. Furthermore, the western Llama models typically outperform Qwen models from China of comparable size. The reasoning chain prompting approach improves overall prediction performance for LLama3-70B, Llama3-8b and Qwen-72B. Notably, Qwen-7B performs worse overall with reasoning prompting. However, while overall weighted F1-score does not improve with reasoning, it significantly alters prediction behavior. Specifically, without reasoning and the wikipedia prompts, Qwen-7B predicts FOR in $95.1\%$ of votes and correctly identifies only $19$ ABSTENTION instances, resulting in an F1-score of just $0.02$ for this class. Introducing reasoning reduces FOR predictions to $83.7\%$ of votes, and crucially, increases the correct identification of ABSTENTION instances to $199$, thereby improving the ABSTENTION F1-score to $0.110$. The AGAINST F1-score also shows an increase from $0.220$ to $0.280$. This demonstrates that for Qwen-7B, the reasoning prompt induces a trade-off: while the aggregate score does not improve, it sacrifices performance on the majority class for a notable increase in sensitivity towards the minority classes.
@@
Interestingly, the type of persona prompts impact models in different ways. While the smaller models LLama3-8B and Qwen-7B can profit from more detailed Wikipedia prompts, the larger models perform better with short attribute prompts.
@@
We note that the best performing approach also correctly predicts the group line, which we define as the majority position of a group per proposal, in $86.32\%$ of cases. When excluding ABSTENTION group lines, which were never predicted as the majority position of a group by the model, $90.16\%$ of the group lines are correctly predicted. The full group names and exact group line predictions compared to the actual group lines are shown in Appendix A. The group lines of center-left and progressive parties is predicted correctly in most cases (Renew, S&D, Greens/EFA), indicating that most prediction mistakes come from politicians that vote against the group line. On the other hand for groups at the edge of the political spectrum (GUE/NGL, ECR, ID) the group line was often inaccurately predicted.
@@
Next, we show the distribution of vote predictions disaggregated by European group in Figure 2. We observe that the model consistently overestimates the number of FOR votes and significantly underestimates the number of ABSTENTION votes. We note that this trend holds across both Llama models. While the Qwen models predict more ABSTENTIONs, they have low precision with the best model achieving $0.129$ and recall ($0.095$) when doing so (cf. Figure 5). All models predict the votes of center-left and progressive groups (S&D, Renew, Greens/EFA) the most accurate and perform worst for groups at the edge of the political spectrum (ID, ECR, GUE/NGL). This aligns with previous research on public opinion ma2024algorithmic.

### v3

@@ Abstract
Large Language Models (LLMs) display remarkable capabilities to understand or even produce political discourse but have been found to consistently exhibit a progressive left-leaning bias. At the same time, so-called persona or identity prompts have been shown to produce LLM behavior that aligns with socioeconomic groups with which the base model is not aligned. In this work, we analyze whether zero-shot persona prompting with limited information can accurately predict individual voting decisions and, by aggregation, accurately predict the positions of European groups on a diverse set of policies. We evaluate whether predictions are stable in response to counterfactual arguments, different persona prompts, and generation methods. Finally, we find that we can simulate the voting behavior of Members of the European Parliament reasonably well, achieving a weighted F1 score of approximately 0.793. Our persona dataset of politicians in the 2024 European Parliament and our code are available at the following url: https://github.com/dess-mannheim/european_parliament_simulation.
@@
Maximilian Kreutner1, Marlene Lutz1, Markus Strohmaier1,2,3 1University of Mannheim 2GESIS - Leibnitz Institute for the Social Sciences, Cologne 3CSH Vienna Correspondence: maximilian.kreutner@uni-mannheim.de
@@ 1 Introduction
Large Language Models (LLMs) are trained on vast amounts of diverse data, encompassing a wide range of attitudes, opinions, and values. Given their broad exposure, it remains unclear to what extent these models inherit various perspectives and how these perspectives might resurface in their responses. As a result, there is a growing body of research dedicated to investigating the latent viewpoints of LLMs, particularly in relation to their political leaning Feng et al. (2023); Wright et al. (2024); Röttger et al. (2024). Studies have found that, by default, LLMs tend to exhibit progressive, left-center political positions Rutinowski et al. (2024); Taubenfeld et al. (2024); Hartmann et al. (2023); Rozado (2023). At the same time, researchers have demonstrated that when properly adapted (e.g., by fine-tuning or prompting), LLMs can effectively understand Bernardelle et al. (2025); Heseltine and Clemm von Hohenberg (2024); Wu et al. (2023) and reproduce Bachmann et al. (2025); Chalkidis and Brandl (2024) the positions of various political groups.
Research objective: Building on these early demonstrations, our work systematically explores the extent to which LLMs can represent the political votes of individual politicians. Specifically, we aim to simulate the voting behavior of all Members of the European Parliament (MEPs) in 2024 by leveraging the ability of LLMs to role-play as different personas. This type of persona prompting has been shown to allow LLMs to make predictions that better align with the points of view of various sociodemographic groups in certain contexts Argyle et al. (2023); Tseng et al. (2024).
The case of the European Parliament is particularly interesting, as MEPs are often part of both a national party and a European group. Therefore, MEPs’ personal preferences and national party lines can significantly influence the decision to dissent from their corresponding European Group Willumsen (2023). Although individual voting decisions in the early 2000s were still mostly determined by the party lines of national parties Hix (2002), the European Parliament has since undergone a phase of "Empowerment," where majority coalitions are policy-specific, which is unique compared to most national parliaments Hix and Høyland (2013).
@@
Approach: We use personas to prompt four LLMs to adopt the identity of each MEP and investigate the extent to which these models can simulate their voting behavior (cf. Figure 1). This includes predicting the individual decision-making of MEPs, who often balance competing interests with the group line on a given policy. We compare different models, persona prompts, and generation methods. We perform additional ablation studies to determine which attributes are the best predictors for LLMs.
Results: We find that in the majority of cases, persona-driven LLMs can accurately predict individual votes, achieving a weighted F1 score of 0.793; however, they do not predict abstentions well. Additionally, persona-driven LLMs are able to correctly predict group positions on a diverse set of topics while also diverging from them at a rate comparable to real-world politicians. Our ablation study supports current political research, as the national party is the strongest predictor of the final vote.
In persona-driven simulations, we find evidence supporting three key observations. First, more detailed personas can help weaker models generate more accurate predictions. Second, prompting models with open-text reasoning before forcing multiple-choice options steers models more effectively. Third, models struggle to convincingly replicate the views of personas that diverge from their inherent biases.
@@ 2 Related Work
Predicting politicians’ voting behavior is an active research area in both political science and NLP. Early work used word frequency, TF-IDF, fine-tuned embeddings, CNNs, or LSTMs to analyze bills and politician profiles Gerrish and Blei (2011); Kraft et al. (2016); Nay (2017); Bonica (2018); Kornilova et al. (2018); Yang et al. (2021). Recent methods incorporate social media and socioeconomic data with graph-based models for richer context Karimi et al. (2020); Mou et al. (2021); Feng et al. (2022); Mou et al. (2023). Agent Based approaches with Large Language Models were first utilized by Li et al. (2025) and have shown to outperform previous methods. However, existing studies focus solely on the simpler two-party U.S. system. For the European Parliament, only traditional Machine Learning methods, e.g., Random Forest, have been tested and have been shown to perform only slightly better than the majority baseline Guadarrama Rios et al. (2025). In addition, most recent approaches depend on access to a vast amount of past voting information or social media data for each politician, which is not always feasible in real-world scenarios. In contrast, our approach requires only publicly available information about politicians and any form of debate content.
Previous research has consistently shown that various LLMs have a bias toward progressive left-of-center political views by evaluating them on survey questions or different downstream tasks Chalkidis and Brandl (2024); Chalkidis (2024); Taubenfeld et al. (2024); Rutinowski et al. (2024); Batzner et al. (2024); Feng et al. (2023); Hartmann et al. (2023). Furthermore, human feedback-tuned LLMs show an even larger bias compared to base models Potter et al. (2024); Santurkar et al. (2023).
Several studies have demonstrated that LLMs can engage in "role-playing" by adopting specific personas within their contextual framework Tseng et al. (2024); Chen et al. (2024); Hu and Collier (2024); Beck et al. (2024). In fact, they are so proficient at this that responses generated while impersonating politicians and other public figures are often perceived as more authentic than those given by the individuals themselves Herbold et al. (2024). This role-playing capability has numerous applications. For instance, by using "silicon samples" and personas modeled with various socioeconomic attributes, LLMs can effectively replicate the viewpoints of demographically diverse subpopulations or predict election and voting behavior Ma et al. (2025); von der Heyde et al. (2025); Argyle et al. (2023); Yu et al. (2024); Zhang et al. (2024); Bradshaw et al. (2024); Yang et al. (2024b); Jiang et al. (2024b). Additionally, generating agents based on long form interviews enables LLMs to reproduce participants’ responses on the General Social Survey with $85\%$ accuracy, compared to their own answers given two weeks later Park et al. (2024). We aim to use this capability to predict voting behavior across European groups, particularly those with whom LLMs are less aligned.
@@ 3 Data Collection
We collect vote information from HowTheyVote [^1] https://howtheyvote.eu/, a project dedicated to making European Parliament results transparently accessible to the public. Our focus is on roll-call votes, the only type of vote in the European Parliament where the individual votes of all MEPs are recorded and published. We identify and select the most significant decisions of the 2024 parliament by checking if there is a press release and a debate about the vote. Our final dataset captures 27,770 individual voting decisions from 710 politicians across 47 key votes, which are displayed in Table 9. Three votes are resolutions initiated by the European Parliament itself, instead of proposals by the Committee. For the sake of simplicity, from this point on, we will still refer to all of these votes as proposals. Each MEP can vote in favor of (FOR), against (AGAINST), abstain (ABSTENTION) from a proposal, or not vote at all. While active abstentions are often strategic decisions Mühlböck and Yordanova (2017), the reasons for not voting are not recorded and can vary, such as illnesses or other responsibilities. As a result, we focus solely on predicting votes categorized as FOR, AGAINST, or ABSTENTION.
@@ 4.1 Models
To prevent data leakage, we choose a model with a training cut-off prior to the earliest vote in our dataset on 16.01.2024: Llama3 Dubey et al. (2024). In addition, we chose a second model that was not developed in the west with comparable performance on the Open LLM Leaderboard Fourrier et al. (2024): Qwen2.5555Qwen2.5’s pretraining cutoff date is not officially disclosed. We considered the risk of data leakage but found it unlikely to be a confounding factor, as the results are comparably worse than LLama3. Yang et al. (2024a). More specifically, we use the models:
@@
Additionally, we experimented with Mixtral Jiang et al. (2024a) as a European model but observed that it did not adhere well to our prompt format, as described in Appendix H. Therefore, we focus on LLama and Qwen for the main section of our paper.
@@ 4.2 Prompt Construction
Persona Construction. To create persona descriptions for each MEP, we consider two strategies. First, we create a combination of demographic and political affiliation attributes, including the full name, gender, age at the time of voting, birthplace, the country they represent, the European group, and the national party they are affiliated with at the time of voting. This approach allows us to precisely control the persona information that the LLM is provided with and enables us to estimate the influence of individual attributes on voting behavior. Second, we use Llama3-70B to summarize the English Wikipedia article associated with each MEP. We investigate whether providing the LLM with longer contextual persona information can improve voting accuracy, similar to the approach used by Park et al. (2024), who used long form interview transcripts to simulate personas.
Proposal Information. It is challenging to provide a short and unbiased description of proposals. The name of the proposal often lacks sufficient detail; for example, one proposal in our dataset is simply titled "Driving licenses". We considered using the official document for each proposal but found that the consolidated legislative document is too long for the context window of LLMs. On the other hand, official press releases, news reports, and summaries provided by the European Committee typically focus only on the positive aspects of the proposals, which could potentially bias the LLMs to always vote in favor of the proposals. Instead, we adopt the theory that a common group position is established by the group’s policy experts and then later adopted by other MEPs Ringe (2009). We choose to use the speeches of each group representative, who are allocated speaking time before each vote to present the group’s position on the respective proposal. The speeches are presented to the LLM in a randomized order, without disclosing the speaker’s identity or group affiliation. Names of politicians, GROUP, and parties appearing in speeches were anonymized by substituting them with placeholders (e.g., [GROUP]; cf. Appendix G.1). We aim to provide a balanced overview of the pro- and contra-arguments regarding the implementation of the proposal while avoiding directly revealing each group’s position. We note that, on average, each speech is nearly 1,500 characters long and constitutes the majority of the prompt.
Response Format. We instruct the model to respond with exactly one of the options: FOR, AGAINST, or ABSTENTION. We further compare two different prompting strategies; reasoning (r), where the model responds first with an open text reasoning chain before choosing one of the options, and no reasoning (nr), where the model answers with one of the options directly. We prompt the model three times for each persona. In all runs, we use a temperature of $0.6$. Our full prompts can be found in Appendix I.
@@ 5 Results
Since the distribution of ground truth votes is heavily imbalanced, with approximately $77\%$ FOR votes, $17\%$ AGAINST votes, and $6\%$ ABSTENTION votes, we decide on the mean weighted F1 score across three runs as our evaluation measure. We report that the responses of all models are robust; i.e., in $87.1\%$ of cases, the models predict the same vote for each persona across all three runs, leading to a low standard deviation of <0.002<0.002 across all runs. We analyze performance across three categories:
@@ 5.1 Persona Results
We show our main results in Table 1. The highest weighted F1-score of $0.793$ is achieved by Llama3-70B when employing reasoning with attribute-based persona prompts. Consistent with expectations, larger models surpass their smaller counterparts in performance. Furthermore, the western Llama models typically outperform Qwen models from China of comparable size. The reasoning chain prompting approach improves overall prediction performance for LLama3-70B, Llama3-8b, and Qwen-72B. Notably, Qwen-7B performs worse overall with reasoning prompting. However, while the overall weighted F1-score does not improve with reasoning, it significantly alters prediction behavior. Specifically, without reasoning and the wikipedia prompts, Qwen-7B predicts FOR in $95.1\%$ of votes and correctly identifies only $19$ ABSTENTION instances, resulting in an F1-score of just $0.02$ for this class. Introducing reasoning reduces FOR predictions to $83.7\%$ of votes and, crucially, increases the correct identification of ABSTENTION instances to $199$, thereby improving the ABSTENTION F1-score to $0.110$. The AGAINST F1-score also shows an increase from $0.220$ to $0.280$. This demonstrates that, for Qwen-7B, the reasoning prompt induces a trade-off: while the aggregate score does not improve, it sacrifices performance on the majority class for a notable increase in sensitivity towards the minority classes.
@@
Interestingly, the type of persona prompts impacts models in different ways. While the smaller models LLama3-8B and Qwen-7B can benefit from more detailed Wikipedia prompts, the larger models perform better with short attribute prompts.
@@
We note that the best performing approach also correctly predicts the group line, which we define as the majority position of a group per proposal, in $86.32\%$ of cases. When excluding ABSTENTION group lines, which were never predicted as the majority position of a group by the model, $90.16\%$ of the group lines are correctly predicted. The full group names and exact group line predictions compared to the actual group lines are shown in Appendix A. The group lines of center-left and progressive parties are predicted correctly in most cases (Renew, S&D, Greens/EFA), indicating that most prediction mistakes come from politicians who vote against the group line. On the other hand, for groups at the edge of the political spectrum (GUE/NGL, ECR, ID), the group line was often inaccurately predicted.
@@
Next, we show the distribution of vote predictions disaggregated by European group in Figure 2. We observe that the model consistently overestimates the number of FOR votes and significantly underestimates the number of ABSTENTION votes. We note that this trend holds across both Llama models. While the Qwen models predict more ABSTENTIONs, they have low precision, with the best model achieving $0.129$ and recall ($0.095$) when doing so (cf. Figure 5). All models predict the votes of center-left and progressive groups (S&D, Renew, Greens/EFA) the most accurately and perform worst for groups at the edge of the political spectrum (ID, ECR, GUE/NGL). This aligns with previous research on public opinion Ma et al. (2025).